# SMReferences

### v1

@@ Abstract
We investigate the collective oscillations of atomic Bose-Einstein condensates (BECs) containing spin-superflow turbulence and observe an enhancement in damping due to the turbulence. As steady-state turbulence is maintained in a spin-1 BEC using a spin-driving technique, we measure the damping rate of the quadrupole oscillation of the BEC across a range of temperatures and show that it exceeds the thermal damping rate predicted by a Landau damping model for BECs without turbulence. To explain the additional damping, we introduce an effective viscosity $\nu_{\text{T}}$, analogous to the turbulent viscosity in classical turbulence. Our measurements suggest that $\nu_{\text{T}}\sim 0.05\times\frac{h}{m}$, where $h$ is Planck’s constant and $m$ is the atomic mass.
@@
An interesting question is whether the notion of turbulent viscosity can be extended to superfluids, which have zero intrinsic viscosity and behave differently from classical fluids due to quantum effects. Although inviscid, superfluids can dissipate kinetic energy via the dynamics of quantum vortices and their interactions with the coexisting normal fluid component [6, 7, 8, 9, 10]. Energy cascades and Kolmogorov scaling behaviors have been observed in turbulent superfluids [11, 12], indicating a remarkable similarity to classical fluids. Previously, the concept of eddy viscosity was adopted in work of turbulence in superfluid helium to account for the excess pressure gradient observed under heat counterflow conditions [13, 14], but the interpretation remains inconclusive due to uncertainties regarding the homogeneity of the turbulence and ambiguities in the normal fluid state [15]. Recent piston-driven shock experiments in atomic Bose-Einstein condensates (BECs) recalled eddy viscosity to explain dissipative shock dynamics [16]. A coherent understanding of turbulent viscosity in superfluid turbulence at the macroscopic level is highly desirable and may have implications for astrophysics, particularly in the context of neutron stars with superfluid interiors [18, 17].
@@
Figure 1: Collective oscillations of a turbulent Bose-Einstein condensate (BEC). (a) Schematic of the experiment. A BEC with internal turbulent flow (gray arrows) is confined in a harmonic potential and undergoes shape oscillations that are damped by the thermal background gas. The turbulence is sustained by resonant rf spin driving, which steadily generates an irregular spin texture (color pattern). (b) Energy flow diagram illustrating two pathways for dissipation of collective excitation energy: direct interaction with thermal components and energy transfer into the internal turbulence. (c) Image of a turbulent BEC after an 18-ms time-of-flight.
In this paper, we present an experimental investigation of the viscous effects of turbulence in atomic BECs. We generate a spin-1 BEC in a non-equilibrium steady state characterized with a turbulent spin-superflow [3, 7, 4, 22] and examine the impact of turbulence on the damping of collective quadrupole oscillations of the BEC. By comparing systems with and without turbulence, and measuring the temperature dependence of the damping rates, we reveal that turbulent BECs exhibit enhanced damping beyond the theoretical predictions for thermal Landau damping [23, 24]. Using a classical turbulence framework [25, 12], we introduce a turbulent kinematic viscosity, $\nu_{\text{T}}$ , to explain the additional damping observed, and our results suggest that $\nu_{\text{T}}\sim 0.05\times h/m$ with $h$ being Planck’s constant and $m$ the atomic mass. This work opens new avenues for the experimental studies of the hydrodynamic properties of superfluid turbulence, using atomic BECs with steady-state turbulence.
@@
Our superfluid system is a BEC of 23Na atoms in the $F$=1 hyperfine state, which has internal spin degrees of freedom [8, 2]. A BEC is prepared initially in the $m_{F}=-1$ state in an optical dipole trap under a uniform external magnetic field. Turbulence is generated using a spin-driving technique [3, 7], where a radio-frequency (RF) magnetic field is applied with an oscillating frequency set to the Lamor frequency. Due to chaotic spin dynamics under resonant spin driving [4, 5, 6], a non-equilibrium steady state emerges with an irregular spin texture, leading to turbulent superflow in the BEC system [Fig. 1(c)]. We emphasize that the driving occurs at the spin energy scale ($\varepsilon_{s}/h\sim 50$ Hz), negligibly small compared to the chemical potential ($\mu/h\sim 1$ kHz) such that the turbulent state is sustained under driving without significant heating. The lifetime of the sample in the non-equilibrium steady state exceeds 30 s, which allows an investigation of the collective oscillations of the turbulent BEC [31, 32].
Collective oscillations of a trapped BEC typically undergo damping in the presence of a coexisting thermal cloud. Thermal dissipation is well described by the Landau damping process, where the energy and momentum of collective excitations are absorbed by thermally excited particles [23, 24]. For a BEC with turbulence, an additional damping mechanism can arise from the nonlinear mixing effect induced by turbulence [Fig. 1(b)], which diffuses the long-wavelength momentum of the collective excitations within the BEC. Energy transport in momentum space has been demonstrated in a recent experiment with turbulent atomic BECs [12].
@@
We excite shape the oscillations of the BEC by periodically modulating the depth of the optical trap as $U(t)=U_{\text{f}}[1+\epsilon\sin(\omega_{d}t)]$ with frequency $\omega_{d}$ and relative amplitude $\epsilon$. For our highly oblate sample with trapping frequencies $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(7.5,15.3,730)$ Hz, two distinct quadrupole modes are identified with oscillation frequencies of about 11 Hz and 25 Hz, respectively. The high (low) frequency mode exhibits in-phase (out-of-phase) oscillations in condensate widths along the $x$ and $y$ directions [9, 10]. In the present work, we investigate a high-frequency mode in which the BEC experiences a higher shear flow during oscillations, thus offering a favorable setting for exploring turbulent viscosity [35].
Figure 2: Responses of BECs to the periodic modulations of the trapping potential. (a) Amplitude magnification factor $\tilde{A}$ and (b) relative phases $\phi$ of the shape oscillations of the driven BECs as functions of the driving frequency $\omega_{d}$ for three different samples: ordinary single-component samples with thermal fractions $\zeta_{\text{th}}=0.3$ (blue) and $=0.6$ (cyan), and a turbulent BEC sample with $\zeta_{\text{th}}=0.6$ (orange). The solid represent indicate Lorentzian curves in (a) and arctangent functions with offsets in (b) fitted to the corresponding data sets (see text). The dashed vertical lines indicate the resonant frequency of the quadrupole oscillation, estimated from the trapping frequencies. The insets show the time evolution of the normalized condensate width $\tilde{W}$ in the sample with $\zeta_{\text{th}}=0.3$ for three different values of $\omega_{d}$. The red lines denote guides for oscillations in phase with the trap modulations.
We characterize the collective excitation mode by measuring the steady-state response of the BEC to trap modulations as a function of the modulation frequency $\omega_{d}$. After a time of $t_{0}=50\frac{2\pi}{\omega_{d}}$, ensuring that the BEC oscillations reach a quasi-steady state, the normalized width $\tilde{W}(t)$ of the BEC along the $y$ direction is tracked by taking a time-of-flight image of the sample for various driving times $t$. Here, $\tilde{W}=W/W_{\text{eq}}$ , with $W$ being the measured width and $W_{\text{eq}}$ the width of an unmodulated BEC for the same atom number $N_{c}$. The relative amplitude $A$ and phase $\phi$ of the BEC oscillations are determined from the fit of $\tilde{W}(t)=1+\alpha_{\tau}A\sin(\omega_{d}t+\phi+\varphi_{\tau})$ to the data, where $\alpha_{\tau}=\sqrt{1+(\omega_{d}\tau)^{2}}$ and $\varphi_{\tau}=\tan^{-1}(\omega_{d}\tau)$ account for the effects of amplitude amplification and the phase change, respectively, due to time-of-flight expansion ($\tau=18$ ms) [35]. In the measurement, the magnitude of the trap modulations was adjusted to hold $A$ to less than 10%, resulting in $\epsilon\leq 1.5\%$. Because $W_{\text{eq}}\propto U_{\text{f}}^{-1/5}$ in a harmonic potential, the change in width at equilibrium is $\Delta W_{\text{eq}}=\frac{1}{5}\epsilon W_{\text{eq}}$ for a small $\epsilon$ , and we estimate the amplitude magnification factor for the driven oscillations as $\tilde{A}=5A/\epsilon$.
@@
In Fig. 2, we present the response spectra $\tilde{A}(\omega_{d})$ and $\phi(\omega_{d})$ for three different samples, in this case BECs with the thermal fraction $\zeta_{\text{th}}=0.32(1)$ (blue markers) and 0.59(1) (light blue), respectively, prepared without RF spin driving, and a turbulent BEC with $\zeta_{\text{th}}=0.57(1)$ (orange). Resonance behavior is evident in $\tilde{A}$, accompanied by a phase change of $\pi$ in $\phi$ as $\omega_{d}$ increases. The resonance frequency $\omega_{\nu}$ and the spectral width $\Gamma$ are determined from the simultaneous fit of $\tilde{A}(\omega_{d})=\frac{A^{\prime}\omega_{\nu}^{2}}{\sqrt{(\omega_{d}^{2}-\omega_{\nu}^{2})^{2}+4(\omega_{d}\Gamma)^{2}}}$ and $\phi(\omega_{d})=\frac{\pi}{2}+\tan^{-1}(\frac{\omega_{d}^{2}-\omega_{\nu}^{2}}{2\Gamma\omega_{d}})-\phi_{0}$ to the measurement data with $A^{\prime}$ and $\phi_{0}$ being free parameters [36]. For ordinary BECs without turbulence, we obtain $\Gamma/2\pi=0.40(3)$ Hz and $\Delta\omega_{\nu}/2\pi=(\omega_{\nu}-\omega_{\nu,0})/2\pi=0.21(3)$ Hz for a low $\zeta_{\text{th}}$, and $\Gamma/2\pi=0.53(5)$ Hz and $\Delta\omega_{\nu}/2\pi=0.44(1)$ Hz for a high $\zeta_{\text{th}}$, where $\omega_{\nu,0}/2\pi=25.3$ Hz is the resonance frequency at zero temperature, calculated from the trap frequencies [35]. As $\zeta_{\text{th}}$ increases, the resonance frequency shifts upward and the damping increases, consistent with previous studies of finite temperature effects [37, 38, 39, 40].
The turbulent BEC exhibits more pronounced spectral broadening and a larger shift with $\Gamma/2\pi=1.1(1)$ Hz and $\Delta\omega_{\nu}/2\pi=1.6(3)$ Hz. The spectral width is more than twice that of an ordinary BEC with a comparable thermal fraction, suggesting that turbulence leads to enhanced damping. However, it is important to note that with RF driving both the thermal cloud and the turbulent BEC consist of three spin components with equal populations [3], resulting in a thermal temperature distinct from that of a single-component system having the same thermal fraction. Thus, to validate the damping enhancement by turbulence, a more comprehensive analysis of the thermal damping effects is needed.
According to a theoretical description of Landau damping for low-lying collective excitations of a trapped BEC [24, 41], the damping rate $\Gamma$ at temperature $k_{\text{B}}T>\mu$ is given by
$\frac{\Gamma}{\omega_{\nu}}=A_{\nu}\sqrt{n_{0}a^{3}}~{}\frac{k_{\text{B}}T}{\mu},$ (1)
where $n_{0}$ is the maximum number density of the condensate, $a$ is the scattering length, and $k_{\text{B}}$ is the Boltzmann constant. The factor $A_{\nu}$ is a dimensionless parameter determined by the spatial geometry of the trap and the characteristics of the collective mode $\nu$. When the number of spin components, $S$, in the thermal cloud increases, the population of thermal particles increases proportionally to $S$, while maintaining an identical momentum distribution at temperature $T$. Therefore, it is reasonable to expect that the damping rate would increase by the same factor as $S$ [8, 42], assuming that the scattering property of the BEC remains the same across different spin components. We propose that $A_{\nu}=SA_{\nu}^{(0)}$, with $A_{\nu}^{(0)}$ representing the value for a single-component system.
We investigate the temperature dependence of the damping rates for BECs with and without turbulence. The temperature is controlled by adjusting the final trap depth of evaporation cooling, where the trapping frequencies of the final trap range from $2\pi\times(4.4,8.9,420)$ Hz to $2\pi\times(7.9,16.3,780)$ Hz, and correspondingly the thermal fraction changes approximately from 0.1 to 0.5 without turbulence and from 0.25 to 0.7 with turbulence. In order to avoid any issues arising from the relative oscillations of the thermal gas coexisting with the BECs [37, 38, 39, 43, 44], we measure the damping rate of the freely decaying oscillations of the BEC after a brief period of trap modulations this time. The modulation period is set to $t_{0}=10\frac{2\pi}{\omega_{d}}$ and the driving frequency is tuned to the red side of the resonance, further away from the closest thermal gas oscillating mode at $2\omega_{y}$. The damping rate $\Gamma$ and the oscillation frequency $\omega_{\nu}$ are determined from a damped sinusoidal function of $\tilde{W}(t)=1+Ae^{-\Gamma t}\sin(\omega_{\nu}t+\phi)$ that fits the experimental data [Fig. 3(b)]. Furthermore, to verify the dependence of $A_{\nu}$ on $S$, we conduct a parallel experiment with atomic samples having two spin components ($S=2$[35], which are equal mixtures of the $m_{F}=\pm 1$ components that are miscible [45, 1]. The field gradient is held below 1 mG/cm on the $xy$ plane, preventing spin dragging effects caused by spatial separation of the spin components [47].
In Fig. 3(c), the relative damping rates $\Gamma/\omega_{\nu}$ are shown for $S=1,2,3$, plotted against the reduced temperature $\widetilde{T}\equiv k_{\text{B}}T\sqrt{n_{0}a^{3}}/\mu$ [48]. The scattering length $a$ varies for different spin components by up to 7% [49] but during the calculation of $\tilde{T}$($\propto a^{4/5}$), we neglect such variations and use the value of $a$ for the $S=1$ sample in the $m_{F}=-1$ state. A linear relationship between $\Gamma/\omega_{\nu}$ and $\widetilde{T}$ is observed as anticipated from Eq. (1). We determine the values of $A_{\nu}$ by applying linear fits to the data sets, resulting in $A_{\nu}=2.1(1)$, $3.8(2)$, and $8.1(4)$ for $S=1,2,3$, respectively. The value of $A_{\nu}$ for $S=1$, that is, $A_{\nu}^{(0)}$, is more than two times smaller than the expected values from theory and experiment [37, 24], which might be due to the dimensional crossover effect for the tight confinement along the $z$ direction [40, 51]. Notably, the value of $A_{\nu}$ for samples with two spin components is found to be consistent with that of $2A_{\nu}^{(0)}$ within the corresponding uncertainty range. This supports our assertion that the thermal damping rate is linearly proportional to the number of spin components.
In light of this finding, the observation that $A_{\nu}>3A_{\nu}^{(0)}$ for turbulent BECs ($S=3$) indicates that the BEC experiences enhanced damping above the expected thermal damping. Thus, we identify an excess damping as $\Gamma_{\text{T}}=\Gamma-3A_{\nu}^{(0)}\tilde{T}\omega_{\nu}$, which we attribute to additional energy dissipation due to the presence of turbulence. The average value of $\Gamma_{\text{T}}$ for our measurements over temperatures is approximately $2\pi\times 0.2$ Hz. In the inset of Fig. 3(c), we plot $\hbar\Gamma_{\text{T}}/\varepsilon_{s}$ with $\varepsilon_{s}$ being the spin interaction energy at the peak atomic density, representing the turbulent kinetic energy scale of the system.
We now present a phenomenological interpretation of the excess damping observed with $\Gamma_{\text{T}}$ using the turbulent viscosity $\nu_{\text{T}}$. In the conventional description of classical turbulence, Reynolds-averaged Navier-Stokes (RANS) equations are widely used, where the time-average effects of turbulence are encapsulated in $\nu_{\text{T}}$, providing a solvable model of the mean flow [1, 11, 12]. Following the reasoning of the RANS, we can introduce the kinematic viscosity $\nu_{\text{T}}$ to our steady-state superfluid turbulence and derive the viscous damping rate $\Gamma_{\text{T}}$ of collective oscillations, associated with $\nu_{\text{T}}$ [25, 13]. For quadrupole oscillations with a mean velocity field of $\bm{v}(\bm{r},t)=(b_{x}x,b_{y}y,b_{z}z)\sin{\omega_{\nu}t}$, the relationship between $\Gamma_{\text{T}}$ and $\nu_{\text{T}}$ is given by [35]
$\Gamma_{\text{T}}=\frac{7\nu_{\text{T}}}{3\bar{R}^{2}}\frac{[3\sum_{\alpha}b_{\alpha}^{2}-(\sum_{\alpha}b_{\alpha})^{2}]}{\sum_{\alpha}b_{\alpha}^{2}(\bar{\omega}/\omega_{\alpha})^{2}}=\frac{7\nu_{\text{T}}}{3\bar{R}^{2}}F(\omega_{\alpha},b_{\alpha}),$ (2)
with $\bar{R}=(R_{x}R_{y}R_{z})^{1/3}$ and $\bar{\omega}=(\omega_{x}\omega_{y}\omega_{z})^{1/3}$, where $R_{\alpha}$ is the Thomas-Fermi (TF) radius of the condensate along the $\alpha$ axis. The dimensionless parameter $F$ characterizes a shear flow in the collective oscillation mode. In Fig. 4, we plot the values of $\nu_{\text{T}}$ estimated from our experimental data using Eq. (2). The average value is obtained as $\nu_{\text{T}}=0.05(4)\times h/m$. Intriguingly, the observed $\nu_{\text{T}}$ is comparable in magnitude to the effective viscosity reported for turbulent superfluid 4He [53, 54, 55], which is associated with energy dissipation within turbulence rather than due to the mean shear flow.
Figure 4: Turbulent kinematic viscosity $\nu_{\text{T}}$, scaled by $\kappa=h/m$, for different $\tilde{T}$. The values of $\nu_{\text{T}}$ were calculated from the excess damping rate $\Gamma_{\text{T}}$ using Eq. (2). The horizontal dashed line indicates the average of the $\nu_{\text{T}}$ values.
A quantitative explanation of the increased damping observed in this study is currently beyond our understanding with regard to the steady-state turbulence generated in spinor BECs under spin driving. A comprehensive description of the energy transfer in this non-equilibrium BEC system is necessary, including the energy spectrum and cascade dynamics within the spin-superflow turbulence [7, 22], along with their alternation through interactions with the surrounding thermal component. This subject is intended for further exploration. Recent studies of plasma turbulence showed that Landau damping can either suppress turbulence [56] or serve as an energy dissipation mechanism to maintain it [57]. The thermal change of the turbulent state could lead to temperature dependence of $\Gamma_{\text{T}}$, most likely implied in Fig. 4.

### v2

# Enhancement of damping in a turbulent atomic Bose-Einstein condensate

@@ Abstract
Turbulence enhances momentum transport in classical fluids, effectively increasing their viscosity. We investigate an analogous effect in a superfluid by measuring the damping of collective oscillations in an atomic Bose-Einstein condensate (BEC) containing stationary spin-superflow turbulence. Using continuous spin driving to maintain turbulence in a spin-1 23Na BEC, we excite its quadrupole mode and measure the damping rate over a range of temperatures. The damping consistently exceeds the Landau-damping rate expected for an equilibrium, non-turbulent BEC. The enhancement likely originates from two complementary processes: direct energy transfer from the mode to turbulent condensate fluctuations and turbulence-induced modification of the thermal cloud that amplifies Landau damping. These results establish collective-mode damping as a sensitive probe of momentum transport in superfluid turbulence.
@@
An intriguing question is whether the notion of turbulent viscosity can be extended to superfluids, which, due to their quantum nature, exhibit zero intrinsic viscosity [6, 7]. Despite their inviscid character, superfluids can still dissipate kinetic energy through the dynamics of quantum vortices [8, 9] and their interactions with the coexisting normal fluid component [10, 11]. Notably, energy cascades and Kolmogorov-like scaling behaviors have been observed in turbulent superfluids [12, 13, 14], pointing to a striking resemblance with classical turbulence. Early studies of superfluid helium turbulence introduced the concept of eddy viscosity to account for excess pressure gradients observed under heat counterflow conditions [15, 16], but the interpretation remained inconclusive due to uncertainties in turbulence homogeneity and ambiguities regarding the state of the normal fluid [17]. More recently, piston-driven shock experiments in atomic Bose-Einstein condensates (BECs) revisited the eddy viscosity framework to describe dissipative shock dynamics [18]. A coherent understanding of turbulence-induced viscous effects in superfluids is therefore highly desirable, with potential implications for astrophysical contexts such as neutron stars, which are believed to host superfluid interiors [20, 19].
@@
Figure 1: Collective oscillations of a turbulent Bose-Einstein condensate (BEC). (a) Schematic of the experiment. A BEC with internal turbulent flow (gray arrows) is confined in a harmonic potential and undergoes shape oscillations. The turbulence is sustained by resonant RF spin driving, which steadily generates an irregular spin texture (color pattern). (b) Energy flow diagram illustrating two pathways for dissipation of collective excitation energy: direct interaction with thermal components and energy transfer into the internal turbulence. (c) Image of a turbulent BEC after an 18-ms time-of-flight.
In this Letter, we experimentally investigate turbulence-enhanced momentum transport in atomic BECs. Using continuous spin driving, we create a spin-1 BEC in a non-equilibrium steady state that hosts turbulent spin-superflow [22, 23, 21]. We then examine how this turbulence affects the damping of collective quadrupole oscillations of the BEC [Fig. 1(a)]. By comparing systems with and without turbulence, and measuring the temperature dependence of the damping rates, we find that turbulent BECs exhibit enhanced damping relative to the Landau-damping expectation for equilibrium condensates [25, 24]. We express this excess in terms of an effective kinematic viscosity, $\nu_{\text{T}}$, directly analogous to the turbulent viscosity of classical fluids [26, 2], and discuss plausible pathways for damping enhancement.
@@
Our superfluid system consists of a BEC of 23Na atoms in the FF=1 hyperfine state, which has internal spin degrees of freedom. We prepare a BEC initially in the $m_{F}=-1$ spin state within an optical dipole trap (ODT) under a uniform external magnetic field. Turbulence is generated using a spin-driving technique as described in [22], in which a radio-frequency (RF) magnetic field is applied at the Larmor resonance frequency. Under resonant spin driving the spin dynamics becomes chaotic [27, 28], giving rise to an irregular spin texture across the BEC [29]. With continuous driving, the system evolves to a non-equilibrium steady state with equal populations of the three, $m_{F}=\pm 1,0$ spin components, where the spin texture is persistently randomized and the associated superflow turbulence is sustained over time [22, 29]. Figure 1(c) shows a time-of-flight image of such a turbulent BEC, where turbulence is evidenced by irregular density modulations. The heating rate from the resonant spin driving is negligible, allowing a turbulent BEC to maintain a long lifetime comparable to the vacuum-limited lifetime (>30>30 s) [22]. This gentle turbulence generation scheme enables the investigation of collective oscillations in a BEC sustained in a steady turbulent state.
Collective oscillations of a trapped BEC typically exhibit damping in the presence of a coexisting thermal cloud. This thermal dissipation is well described by the Landau damping mechanism, in which the energy and momentum of collective excitations are transferred to thermally excited particles [30, 25]. For low-lying collective excitations, the damping rate $\Gamma_{th}$ at temperatures $k_{\text{B}}T>\mu$ is given by [24]
$\frac{\Gamma_{th}}{\omega_{\nu}}=A_{\nu}\sqrt{n_{0}a^{3}}~\frac{k_{\text{B}}T}{\mu},$ (1)
where $\omega_{\nu}$ is the oscillation frequency, $n_{0}$ is the peak condensate density, $a$ is the scattering length, $k_{\text{B}}$ is the Boltzmann constant, and $\mu$ is the chemical potential. The factor $A_{\nu}$ is a dimensionless parameter determined by the trap geometry and the characteristics of the collective mode $\nu$ [31, 32]. This form of thermal damping has been extensively demonstrated in previous experiments for various collective modes of trapped BECs [33, 34, 35, 37, 38, 36].
In a turbulent BEC, nonlinear coupling between collective modes can provide an extra damping channel by diffusing long-wavelength momentum throughout the turbulent medium. Monitoring how turbulence alters the damping behavior of low-lying collective excitations therefore provides a sensitive probe of turbulence-induced dissipation. This approach is particularly effective because it relies on well-established collective modes that have been widely used to characterize finite-temperature effects in BECs.
Figure 2: Observation of damped oscillations of a turbulent BEC. (a) Experimental sequence. The sample temperature is controlled by adjusting the final trap depth $U_{f}$ during evaporation cooling. After cooling, sample is held for a few seconds to thermalize and damp out residual motion. A resonant RF magnetic field is represented as a green solid line, continuously applied to generate and sustain turbulence. The hatched area indicates the period of transition to steady turbulence. After a steady turbulent state is prepared, the trap is perturbatively modulated during a time $t_{0}$, to excite the quadrupole mode. After a variable hold time $t_{h}$, an absorption image is taken after time-of-flight of 18-ms. (b) Time evolution of the normalized condensate width $\widetilde{W}$ along the $y$ direction as a function of the hold time $t_{h}$. Images at top were taken at $t_{h}=36$-ms, $168$-ms, and $299$-ms, from left to right, respectively. Each data point represents a single measurement and the solid red line is a damped sinusoidal function fit to the mean values of the data. This fit yields an oscillation frequency $\omega_{\nu}=18.8(9)$ Hz and damping rate $\Gamma=0.60(9)$ Hz. The thermal fraction of the sample was $0.44(1)$.
@@
In Fig. 2(a), we delineate our experimental procedure to investigate the damping of collective oscillations. After preparing a BEC in a steady turbulent state by applying a RF field, we drive collective excitations by modulating the power of the ODT laser beam. The depth of the optical trap is modulated as $U(t)=U_{f}[1+\epsilon\sin(\omega_{d}t)]$ for a short period set to $t_{0}=10\frac{2\pi}{\omega_{d}}$, with its frequency $\omega_{d}$ and relative amplitude factor $\epsilon$. For our highly oblate sample with trapping frequency ratios of $\omega_{x}:\omega_{y}:\omega_{z}\approx 1:2:100$, two distinct quadrupole modes are identified with oscillation frequencies of $\omega_{\nu}\approx 0.8\,\omega_{y}$ and $1.6\,\omega_{y}$ [39]. The high (low) frequency mode exhibits in-phase (out-of-phase) oscillations in condensate widths along the $x$ and $y$ directions [40, 41]. In the present work, we investigate the high-frequency mode in which the BEC experiences a higher shear flow during oscillations, offering a favorable setting for exploring turbulent viscosity [39, 42, 43]. To selectively excite the quadrupole mode and minimize interference from other modes, $\omega_{d}$ was tuned to the red side of the resonance, further away from the closest thermal gas mode at $2\omega_{y}$ [33, 34, 37, 39].
The oscillations of the BEC are measured by tracking the normalized width $\widetilde{W}(t_{h})$ along the $y$ direction after a variable hold time $t_{h}$ and subsequent time-of-flight imaging [Fig. S2(b)]. Here, $\widetilde{W}=W/W_{\text{eq}}$ , with $W$ being the measured width and $W_{\text{eq}}$ the width of an unmodulated BEC for the measured atom number $N_{c}$ of the condensate [39]. The oscillating behavior is well described by an exponentially damped sinusoid, $\widetilde{W}(t_{h})=1+Ae^{-\Gamma t_{h}}\sin(\omega_{\nu}t_{h}+\phi)$, where $A$ and $\phi$ are the relative amplitude and phase of the oscillation, respectively. The persistence of a well-defined mode indicates that the hydrodynamic nature is preserved in the turbulent BEC, ensuring that the mode’s damping rate offers a quantitative probe of the effect of the turbulence. In our measurements, we keep the in-trap oscillation amplitude below 10% [39].
To demonstrate damping enhancement due to turbulence, we first need an appropriate reference for comparison. The most natural baseline is a thermal equilibrium sample that contains the same numbers of thermal and condensed atoms, $N_{th}$ and $N_{c}$, as the turbulent sample and also has the same spin composition. In the steady turbulent state under continuous spin driving, the thermal cloud is an equal mixture of three spin components [22], and the corresponding equilibrium temperature is $k_{\text{B}}T=0.94\hbar\overline{\omega}({N_{th}}/\mathcal{D}_{s})^{1/3}$ with $\bar{\omega}=(\omega_{x}\omega_{y}\omega_{z})^{1/3}$ and $\mathcal{D}_{s}$=3 denoting the number of spin components [25]. However, a difficulty arises from the fact that a spin-1 BEC at thermal equilibrium tends to develop magnetic ordering due to spin interactions [44], which disfavors a perfectly spin-symmetric thermal cloud. Therefore, directly measuring the reference damping rate for a $\mathcal{D}_{s}$=3 thermal-equilibrium sample is impractical. Nevertheless, since the number of thermal atoms scales with $\mathcal{D}_{s}$ while their momentum distribution at temperature $T$ remains unchanged, the damping rate should increase by the same factor, provided the scattering properties are identical for all spin components [44, 45]. Accordingly, we estimate the reference damping rate from Eq. (1) using $A_{\nu}=\mathcal{D}_{s}A_{\nu}^{(0)}$, where $A_{\nu}^{(0)}$ is the value of a single component system at thermal equilibrium.
With this reference damping rate in hand, we investigate the temperature dependence of the damping rates for both spin-driven turbulent BECs ($\mathcal{D}_{s}=3$) and single-component BECs ($\mathcal{D}_{s}=1$) prepared without RF spin driving. The temperature is controlled by adjusting the final trap depth $U_{f}$ during evaporation cooling [Fig. 2(a)]. As $U_{f}$ increases, the final trapping frequencies range from $2\pi\times(4.4,8.9,420)$ Hz to $2\pi\times(7.9,16.3,780)$ Hz, and correspondingly the thermal fraction changes approximately from 0.25 to 0.7 with turbulence and from 0.1 to 0.5 without turbulence.
@@
Furthermore, to verify the dependence of $A_{\nu}$ on $\mathcal{D}_{s}$, we conduct a parallel experiment with thermal equilibrium samples having two spin components ($\mathcal{D}_{s}=2$) [Fig. 3(a)], which are equal mixtures of the $m_{F}=\pm 1$ components that are miscible for our 23Na BEC system [46]. To prepare them, the atoms are transferred to the $m_{F}=0$ state using a rapid adiabatic passage with an RF magnetic field, followed by a pulse of a strong magnetic field gradient to purify the spin state. We then applied a $\pi/2$ resonant RF magnetic pulse to form an equal mixture of the $m_{F}=\pm 1$ components. After the preparation of the spin mixture, the quadratic Zeeman energy is changed from positive to small negative value to prevent spontaneous formation of the $m_{F}=0$ population and stabilize the two-component system [47, 48]. In all the damping-rate measurements, we nulled the magnetic field gradient to below 0.1 mG/cm to suppress spin-drag effects in the collective oscillations [49]. The residual field gradient was calibrated and minimized using a Ramsey interferometry technique [50, 22, 48].
Figure 3(b) shows the measurement results of the relative damping rates $\Gamma/\omega_{\nu}$ for samples with one, two and three spin components ($\mathcal{D}_{s}=1,2,3$) as functions of the reduced temperature $\widetilde{T}\equiv k_{\text{B}}T\sqrt{n_{0}a^{3}}/\mu$. The scattering length $a$ varies by up to 7% among the spin components [51], but because $\widetilde{T}\propto a^{4/5}$ [52], this variation alters $\widetilde{T}$ by $\lesssim 6$%. For simplicity, we use the $m_{F}=-1$ value of $a$ for all data sets [53]. For $\mathcal{D}_{s}=1,2$, we extract the Landau-damping prefactor $A_{\nu}$ by linear fits to Eq. (1), obtaining $A_{\nu}=2.1(1)$ and $3.8(2)$, respectively. For the turbulent $\mathcal{D}_{s}=3$ sample, we do not assume the applicability of Eq. (1); instead, we use $\widetilde{T}$ only as a scaling variable to facilitate a quantitative comparison with an equilibrated reference, yielding an effective slope $A_{\nu}=8.1(4)$.
We note that the single-component value, $A_{\nu}^{(0)}\equiv A_{\nu}(\mathcal{D}_{s}=1)$, is smaller than typical values reported for cylindrically symmetric traps [33]. Such variations are expected because $A_{\nu}$ depends sensitively on trap geometry and mode structure in a trapped gas [24, 31, 32]. We attribute the reduced $A_{\nu}^{(0)}$ to the highly oblate trap geometry and tight confinement along $z$, which may place the system in a dimensional-crossover regime [38, 55]. The two-component result satisfies $A_{\nu}=2A_{\nu}^{(0)}$ within experimental uncertainty, confirming that the thermal damping rate scales linearly with the number of spin components. We therefore use the measured $A_{\nu}^{(0)}$ as an empirical baseline for our specific trap geometry and quadrupole mode.
Relative to this calibrated benchmark, the turbulent three-component sample yields $A_{\nu}>~3A_{\nu}^{(0)}$, indicating damping beyond the equilibrium expectation. We quantify the excess damping as
$\Gamma_{\text{T}}=\Gamma-3A_{\nu}^{(0)}\widetilde{T}\omega_{\nu}$ (2)
and attribute it to turbulence-induced damping. Averaged over the temperature range explored, we find $\Gamma_{\text{T}}\approx 2\pi\times 0.2$ Hz.
Following the phenomenological approach in classical fluid dynamics, we recast the excess damping $\Gamma_{\text{T}}$ in terms of turbulent viscosity $\nu_{\text{T}}$ [26, 56]. In classical turbulence, Reynolds-averaged Navier-Stokes (RANS) equations absorb the time-averaged influence of velocity fluctuations into a single parameter $\nu_{\text{T}}$, providing a closed model of mean flow [1, 2]. Adopting this framework to our steady spin-superfluid turbulence and modeling the quadrupole oscillations with the mean velocity field of $\bm{v}(\bm{r},t)=(b_{x}x,b_{y}y,b_{z}z)\sin{\omega_{\nu}t}$, we relate the measured excess damping to a shear-stress action proportional to $\nu_{\text{T}}$ [39]. We obtain an effective kinematic viscosity of $\nu_{\text{T}}=0.05(2)\kappa$, where $\kappa=h/m$ with $h$ being the Planck constant and $m$ the atomic mass. Interestingly, this value is of the same order as the effective viscosity $\sim 0.1\kappa$ reported for turbulent superfluid 4He in the $T\rightarrow 0$ limit [57, 59, 58]. However, direct comparison is tentative because the measurements stem from different experimental contexts and no unified theoretical framework yet connects them [60].
The key question now concerns the physical origin of the enhanced damping. Two mechanisms are most plausible. First, the quadrupole mode’s shear stress can transfer energy directly to the turbulent condensate fluctuations, a purely hydrodynamic pathway that underlies our RANS-based estimation of turbulent viscosity. Second, turbulence can amplify Landau damping by reshaping the surrounding thermal cloud. Because Landau damping of low-lying modes is highly sensitive to the occupation of thermal states near the chemical potential, even modest, non-equilibrium turbulence can redistribute those occupations and alter the damping rate. This feedback loop goes beyond the conventional view of the thermal component as merely a high-momentum energy sink via mutual friction with quantum vortices [61]. Instead, it highlights the full two-fluid character of the superfluid system and may have important implications for wave turbulence in the atomic BEC system. Disentangling the relative contributions of these two channels will require a detailed characterization of the spin-superflow turbulence, such as its energy spectrum, cascade dynamics, and reciprocal coupling to the thermal gas [21, 23].